# Unveiling a 36 Billion Solar Mass Black Hole at the Centre of the Cosmic Horseshoe Gravitational Lens

### v1

@@ Appendix A Conjugated regions
Instead of selecting pairs of conjugated points that are expected to correspond to roughly the same location in the source plane, we opted for choosing pairs of conjugated regions that are expected to overlap at the source plane. This approach is particularly suitable when identifying conjugated points is unclear, such as in the case of faint counter-images. The method is implemented as follows:
@@
In Fig. 12, we illustrate the method. The left panel shows a pair of conjugated regions in the image plane, which are presumably part of the lensed source. The central panel displays the corresponding regions in the source plane after being delensed. In this scenario, since the regions do not overlap in the source plane, the lens macro model will be penalised. The right panel shows the case where the regions overlap in the source plane after being delensed. In this scenario, the lens macro model is accepted.
@@ Appendix B Elliptical Power-Law (EPL) mass model
$\epsilon_{1}=\frac{1-q^{\text{lens}}}{1+q^{\text{lens}}}\sin{2\phi^{\text{lens}}},\quad\quad\epsilon_{2}=\frac{1-q^{\text{lens}}}{1+q^{\text{lens}}}\cos{2\phi^{\text{lens}}}.$ (19)
@@
$\epsilon^{\text{sh}}_{1}$ $-0.026^{+0.001}_{-0.001}$
$\epsilon^{\text{sh}}_{2}$ $-0.008^{+0.001}_{-0.001}$
@@ Appendix C Priors
$M^{\text{DM}}_{200}$ $\log_{10}U[10^{10},10^{15}]$ mass at $200$ times the critical density of the Universe $M_{\odot}$
@@ Appendix D Alternative mass models results
Table 7: Posterior median and $1\sigma$ uncertainties for the parameters in the perturbation models.
Parameter M2 M3 M4 M5 M6 M7 M8
$i$ $66^{+14}_{-10}$ $61^{+11}_{-7}$ $67^{+15}_{-12}$ $59^{+5}_{-5}$ $64^{+17}_{-11}$ $66^{+14}_{-11}$ $69^{+13}_{-12}$
$\beta^{0}_{\text{star}}$ $-0.14^{+0.20}_{-0.20}$ $0.04^{+0.34}_{-0.34}$ $0.09^{+0.12}_{-0.15}$ $-0.14^{+0.16}_{-0.14}$ $0.03^{+0.05}_{-0.05}$ $-0.01^{+0.10}_{-0.16}$ $-0.17^{+0.14}_{-0.16}$
@@
$\log_{10}M_{\text{BH}}$ $10.57^{+0.07}_{-0.09}$ $10.45^{+0.11}_{-0.14}$ $10.53^{+0.10}_{-0.11}$ $10.56^{+0.08}_{-0.08}$ $10.55^{+0.08}_{-0.09}$ $10.55^{+0.08}_{-0.08}$ $10.51^{+0.07}_{-0.09}$
$\epsilon^{\text{sh}}_{1}$ $0.02^{+0.01}_{-0.01}$ $0.03^{+0.01}_{-0.01}$ $0.03^{+0.01}_{-0.01}$ $0.03^{+0.01}_{-0.01}$ $0.02^{+0.01}_{-0.01}$ $0.03^{+0.01}_{-0.01}$ $0.02^{+0.02}_{-0.01}$
$\epsilon^{\text{sh}}_{2}$ $0.05^{+0.01}_{-0.01}$ $0.06^{+0.01}_{-0.01}$ $0.05^{+0.01}_{-0.01}$ $0.06^{+0.01}_{-0.01}$ $0.05^{+0.01}_{-0.01}$ $0.05^{+0.02}_{-0.01}$ $0.05^{+0.01}_{-0.01}$
$M_{\text{Ein}}\left[\frac{M_{\sun}}{10^{12}}\right]$ $5.45^{+0.03}_{-0.03}$ $5.45^{+0.02}_{-0.02}$ $5.45^{+0.02}_{-0.02}$ $5.45^{+0.02}_{-0.02}$ $5.46^{+0.02}_{-0.03}$ $5.45^{+0.02}_{-0.02}$
In Table 8 we present the results of the alternative model M9.
Table 8: Inferred median and $1\sigma$ credible intervals for the parameters of model M9
. Parameter Posterior (median with $1\sigma$ uncertainties) $i$ $71^{+10}_{-8}$ $\beta^{0}_{\text{star}}$ $-0.21^{+0.17}_{-0.13}$ $\beta^{1}_{\text{star}}$ $-0.05^{+0.41}_{-0.15}$ $\beta^{2}_{\text{star}}$ $-0.02^{+0.16}_{-0.17}$ $\Upsilon^{0}_{\star}$ $5.17^{+1.19}_{-0.84}$ $\Upsilon^{1}_{\star}$ $3.21^{+0.79}_{-0.48}$ $\Upsilon^{2}_{\star}$ $1.66^{+1.87}_{-0.83}$ $\log_{10}\rho_{s}$ $-2.41^{+0.04}_{-0.05}$ $q_{\text{DM}}$ $0.97^{+0.03}_{-0.02}$ $\gamma_{\text{DM}}$ $1.08^{+0.06}_{-0.07}$ $\log_{10}M_{\text{BH}}$ $10.50^{+0.10}_{-0.32}$ $\epsilon^{\text{sh}}_{1}$ $0.03^{+0.01}_{-0.01}$ $\epsilon^{\text{sh}}_{2}$ $0.06^{+0.01}_{-0.01}$ $M_{\text{Ein}}\left[\frac{M_{\sun}}{10^{12}}\right]$ $5.45^{+0.02}_{-0.03}$
In Table 9 we present the results of the alternative model M10.
Table 9: Inferred median and $1\sigma$ credible intervals for the parameters of model M10
. Parameter Posterior (median with $1\sigma$ uncertainties) $i$ $75^{+10}_{-8}$ $\beta^{0}_{\text{star}}$ $-0.07^{+0.21}_{-0.62}$ $\beta^{1}_{\text{star}}$ $-0.51^{+0.87}_{-0.40}$ $\beta^{2}_{\text{star}}$ $-0.34^{+0.36}_{-0.39}$ $\beta^{3}_{\text{star}}$ $0.12^{+0.18}_{-0.32}$ $\beta^{4}_{\text{star}}$ $0.37^{+0.16}_{-0.42}$ $\beta^{5}_{\text{star}}$ $0.05^{+0.21}_{-0.27}$ $\beta^{6}_{\text{star}}$ $-0.43^{+0.13}_{-0.17}$ $\beta^{7}_{\text{star}}$ $0.13^{+0.09}_{-0.05}$ $\Upsilon_{0}$ $4.38^{+0.57}_{-1.31}$ $\upsilon_{0}$ $0.53^{+0.08}_{-0.08}$ $\delta$ $7.77^{+1.72}_{-7.20}$ $\log_{10}\rho_{s}$ $-2.38^{+0.01}_{-0.01}$ $\log_{10}M_{\text{BH}}$ $10.55^{+0.10}_{-0.07}$ $\epsilon^{\text{sh}}_{1}$ $0.00^{+0.01}_{-0.01}$ $\epsilon^{\text{sh}}_{2}$ $0.05^{+0.01}_{-0.01}$ $M_{\text{Ein}}\left[\frac{M_{\sun}}{10^{12}}\right]$ $5.45^{+0.02}_{-0.02}$
In Table 10 we present the results of the alternative models M11M13, where we used the mass-concentration relation from Ludlow et al. (2016) to set the DM scale radius.
Table 10: Inferred median and $1\sigma$ credible intervals for the parameters of models M11M13. Note that the DM scale radius is not a free parameter in these models, but setted using the mass-concentration relation from Ludlow et al. (2016).
. Parameter M11 M12 M13 $i$ $75^{+8}_{-10}$ $76^{+9}_{-12}$ $65^{+20}_{-11}$ $\beta^{0}_{\text{star}}$ $-0.01^{+0.24}_{-0.24}$ $-0.02^{+0.24}_{-0.28}$ $0.08^{+0.25}_{-0.15}$ $\beta^{1}_{\text{star}}$ $0.42^{+0.06}_{-0.11}$ $0.34^{+0.09}_{-0.10}$ $-0.39^{+0.22}_{-0.07}$ $\beta^{2}_{\text{star}}$ $-0.29^{+0.19}_{-0.14}$ $-0.10^{+0.24}_{-0.23}$ $-0.26^{+0.16}_{-0.11}$ $\Upsilon^{0}_{\star}$ $5.39^{+1.81}_{-1.33}$ $5.79^{+2.40}_{-1.64}$ $4.51^{+1.24}_{-0.89}$ $\Upsilon^{1}_{\star}$ $3.63^{+0.32}_{-0.35}$ $3.14^{+0.16}_{-0.22}$ $3.16^{+0.34}_{-0.20}$ $\Upsilon^{2}_{\star}$ $2.75^{+0.61}_{-1.04}$ $1.55^{+0.70}_{-0.77}$ $2.69^{+0.48}_{-0.62}$ $M^{\text{DM}}_{200}\left[\frac{M_{\sun}}{10^{13}}\right]$ $8.36^{+1.48}_{-1.22}$ $11.1^{+4.36}_{-2.84}$ $9.52^{+1.28}_{-7.92}$ $q_{\text{DM}}$ $0.99^{+0.01}_{-0.01}$ $0.99^{+0.01}_{-0.01}$ $0.99^{+0.01}_{-0.01}$ $r_{s}$ $27.96^{+2.10}_{-1.90}$ $44.89^{+7.14}_{-5.51}$ $20.97^{+1.21}_{-0.80}$ $\gamma_{\text{DM}}$ $1.29^{+0.05}_{-0.04}$ $1.34^{+0.07}_{-0.07}$ $1.12^{+0.03}_{-0.04}$ $\log_{10}M_{\text{BH}}$ $10.15^{+0.17}_{-0.30}$ $10.33^{+0.07}_{-0.13}$ $10.59^{+0.04}_{-0.10}$ $\epsilon^{\text{sh}}_{1}$ $0.03^{+0.01}_{-0.01}$ $0.03^{+0.01}_{-0.01}$ $0.01^{+0.01}_{-0.01}$ $\epsilon^{\text{sh}}_{2}$ $0.06^{+0.01}_{-0.01}$ $0.06^{+0.01}_{-0.01}$ $0.06^{+0.01}_{-0.01}$ $M_{\text{Ein}}\left[\frac{M_{\sun}}{10^{12}}\right]$ $5.44^{+0.02}_{-0.02}$ $5.43^{+0.02}_{-0.02}$ $5.43^{+0.02}_{-0.02}$
In Table 11 we present the results of the alternative models M14 and M15 in which we did not include the SMBH.
@@
Table 11: Inferred median and $1\sigma$ credible intervals for the parameters of models without the SMBH.
. Parameter M14 M15 $i$ $79^{+7}_{-9}$ $79^{+7}_{-10}$ $\beta_{\text{star}}$ $0.28^{+0.04}_{-0.04}$ $0.42^{+0.05}_{-0.08}$ $\Upsilon^{0}_{\star}$ $4.14^{+0.08}_{-0.07}$ $7.57^{+1.71}_{-2.26}$ $\Upsilon^{1}_{\star}$ $3.90^{+0.19}_{-0.13}$ $\Upsilon^{2}_{\star}$ $1.29^{+0.90}_{-0.70}$ $\log_{10}\rho_{s}$ $-2.40^{+0.01}_{-0.01}$ $-2.36^{+0.01}_{-0.02}$ $q_{\text{DM}}$ $0.99^{+0.01}_{-0.01}$ $0.98^{+0.01}_{-0.01}$ $\epsilon^{\text{sh}}_{1}$ $0.02^{+0.01}_{-0.01}$ $0.02^{+0.01}_{-0.01}$ $\epsilon^{\text{sh}}_{2}$ $0.05^{+0.01}_{-0.01}$ $0.05^{+0.01}_{-0.01}$ $M_{\text{Ein}}\left[\frac{M_{\sun}}{10^{12}}\right]$ $5.45^{+0.02}_{-0.03}$ $5.45^{+0.03}_{-0.03}$

### v3

@@ Appendix A Conjugated regions
Instead of selecting pairs of conjugated points in the lens plane that are expected to map to the same location in the source plane, we selected pairs of conjugated regions in the lens plane that are expected to overlap when mapped back to the source plane. This approach is particularly suitable when identifying conjugated points is unclear, such as in the case of faint counter-images. The method is implemented as follows:
@@
In Fig. 12, we illustrate the method. The left panel shows a pair of conjugated regions in the image plane, which are presumably part of the lensed source. The central panel displays the corresponding regions in the source plane after being delensed. In this scenario, since the regions do not overlap in the source plane, the lens macro model will be penalised. The right panel shows the case where the regions overlap in the source plane after being delensed, and the lens macro model is accepted.
@@ Appendix B Elliptical Power-Law (EPL) mass model
$\epsilon_{1}=\frac{1-q^{\text{lens}}}{1+q^{\text{lens}}}\sin{2\phi^{\text{lens}}},\quad\quad\epsilon_{2}=\frac{1-q^{\text{lens}}}{1+q^{\text{lens}}}\cos{2\phi^{\text{lens}}},$ (19)
which prevents sampling issues related to the periodicity of $\phi^{\text{lens}}$ and the discontinuity at $q^{\text{lens}}=0$ (e.g., Galan et al., 2022; Etherington et al., 2022).
@@
$\epsilon^{\text{sh}}_{1}$ $0.026^{+0.001}_{-0.001}$
$\epsilon^{\text{sh}}_{2}$ $0.008^{+0.001}_{-0.001}$
@@ Appendix C Priors
        $M^{\text{DM}}_{200}$         $\log_{10}U[10^{10},10^{15}]$                  mass at $200$ times the critical         density of the Universe         $10^{13}M_{\sun}$
@@
        $M_{\text{Ein}}$         $N[5.46,0.27]$         Einstein mass within the tangential ring         $10^{12}M_{\sun}$
@@ Appendix D Alternative mass models results
Table 7: Posterior median and $1\sigma$ uncertainties for the parameters in the fiducial and perturbation models. Parameter units are given in Table 6.
Parameter M1 M2 M3 M4 M5 M6 M7 M8
$i$ $65^{+15}_{-11}$ $66^{+14}_{-10}$ $61^{+11}_{-7}$ $67^{+15}_{-12}$ $59^{+5}_{-5}$ $64^{+17}_{-11}$ $66^{+14}_{-11}$ $69^{+13}_{-12}$
$\beta^{0}_{\text{star}}$ $0.07^{+0.06}_{-0.10}$ $-0.14^{+0.20}_{-0.20}$ $0.04^{+0.34}_{-0.34}$ $0.09^{+0.12}_{-0.15}$ $-0.14^{+0.16}_{-0.14}$ $0.03^{+0.05}_{-0.05}$ $-0.01^{+0.10}_{-0.16}$ $-0.17^{+0.14}_{-0.16}$
@@
$\log_{10}M_{\text{BH}}$ $10.56^{+0.07}_{-0.08}$ $10.57^{+0.07}_{-0.09}$ $10.45^{+0.11}_{-0.14}$ $10.53^{+0.10}_{-0.11}$ $10.56^{+0.08}_{-0.08}$ $10.55^{+0.08}_{-0.09}$ $10.55^{+0.08}_{-0.08}$ $10.51^{+0.07}_{-0.09}$
$\epsilon^{\text{sh}}_{1}$ $-0.01^{+0.01}_{-0.01}$ $-0.02^{+0.01}_{-0.01}$ $-0.03^{+0.01}_{-0.01}$ $-0.03^{+0.01}_{-0.01}$ $-0.03^{+0.01}_{-0.01}$ $-0.02^{+0.01}_{-0.01}$ $-0.03^{+0.01}_{-0.01}$ $-0.02^{+0.02}_{-0.01}$
$\epsilon^{\text{sh}}_{2}$ $-0.05^{+0.01}_{-0.01}$ $-0.05^{+0.01}_{-0.01}$ $-0.06^{+0.01}_{-0.01}$ $-0.05^{+0.01}_{-0.01}$ $-0.06^{+0.01}_{-0.01}$ $-0.05^{+0.01}_{-0.01}$ $-0.05^{+0.02}_{-0.01}$ $-0.05^{+0.01}_{-0.01}$
$M_{\text{Ein}}$ $5.45^{+0.02}_{-0.03}$ $5.45^{+0.03}_{-0.03}$ $5.45^{+0.02}_{-0.02}$ $5.45^{+0.02}_{-0.02}$ $5.45^{+0.02}_{-0.02}$ $5.46^{+0.02}_{-0.03}$ $5.45^{+0.02}_{-0.02}$
The median and uncertainties of the parameters in the alternative models are summarised in Table 8. Units and priors can be checked in Table 6.
@@
Table 8: Inferred median and $1\sigma$ credible intervals for the parameters of alternative models M9-M13, and for the models without the SMBH contribution (M14 and M15). Parameter units are given in Table 6.
Parameter M9 M10 M11 M12 M13 M14 M15
$i$ $71^{+10}_{-8}$ $75^{+10}_{-8}$ $75^{+8}_{-10}$ $76^{+9}_{-12}$ $65^{+20}_{-11}$ $79^{+7}_{-9}$ $79^{+7}_{-10}$
$\beta^{0}_{\text{star}}$ $-0.21^{+0.17}_{-0.13}$ $-0.07^{+0.21}_{-0.62}$ $-0.01^{+0.24}_{-0.24}$ $-0.02^{+0.24}_{-0.28}$ $0.08^{+0.25}_{-0.15}$ $0.28^{+0.04}_{-0.04}$ $0.42^{+0.05}_{-0.08}$
$\beta^{1}_{\text{star}}$ $-0.05^{+0.41}_{-0.15}$ $-0.51^{+0.87}_{-0.40}$ $0.42^{+0.06}_{-0.11}$ $0.34^{+0.09}_{-0.10}$ $-0.39^{+0.22}_{-0.07}$
$\beta^{2}_{\text{star}}$ $-0.02^{+0.16}_{-0.17}$ $-0.34^{+0.36}_{-0.39}$ $-0.29^{+0.19}_{-0.14}$ $-0.10^{+0.24}_{-0.23}$ $-0.26^{+0.16}_{-0.11}$
$\beta^{3}_{\text{star}}$ $0.12^{+0.18}_{-0.32}$
$\beta^{4}_{\text{star}}$ $0.37^{+0.16}_{-0.42}$
$\beta^{5}_{\text{star}}$ $0.05^{+0.21}_{-0.27}$
$\beta^{6}_{\text{star}}$ $-0.43^{+0.13}_{-0.17}$
$\beta^{7}_{\text{star}}$ $0.13^{+0.09}_{-0.05}$
$\Upsilon^{0}_{\star}$ $5.17^{+1.19}_{-0.84}$ $5.39^{+1.81}_{-1.33}$ $5.79^{+2.40}_{-1.64}$ $4.51^{+1.24}_{-0.89}$ $4.14^{+0.08}_{-0.07}$ $7.57^{+1.71}_{-2.26}$
$\Upsilon^{1}_{\star}$ $3.21^{+0.79}_{-0.48}$ $3.63^{+0.32}_{-0.35}$ $3.14^{+0.16}_{-0.22}$ $3.16^{+0.34}_{-0.20}$ $3.90^{+0.19}_{-0.13}$
$\Upsilon^{2}_{\star}$ $1.66^{+1.87}_{-0.83}$ $2.75^{+0.61}_{-1.04}$ $1.55^{+0.70}_{-0.77}$ $2.69^{+0.48}_{-0.62}$ $1.29^{+0.90}_{-0.70}$
$\Upsilon_{0}$ $4.38^{+0.57}_{-1.31}$
$\upsilon_{0}$ $0.53^{+0.08}_{-0.08}$
$\delta$ $7.77^{+1.72}_{-7.20}$
$\log_{10}\rho_{s}$ $-2.41^{+0.04}_{-0.05}$ $-2.38^{+0.01}_{-0.01}$ $-2.40^{+0.01}_{-0.01}$ $-2.36^{+0.01}_{-0.02}$
$M^{\text{DM}}_{200}$ $8.36^{+1.48}_{-1.22}$ $11.1^{+4.36}_{-2.84}$ $9.52^{+1.28}_{-7.92}$
$q_{\text{DM}}$ $0.97^{+0.03}_{-0.02}$ $0.99^{+0.01}_{-0.01}$ $0.99^{+0.01}_{-0.01}$ $0.99^{+0.01}_{-0.01}$ $0.99^{+0.01}_{-0.01}$ $0.98^{+0.01}_{-0.01}$
$r_{s}$ $27.96^{+2.10}_{-1.90}$ $44.89^{+7.14}_{-5.51}$ $20.97^{+1.21}_{-0.80}$
$\gamma_{\text{DM}}$ $1.08^{+0.06}_{-0.07}$ $1.29^{+0.05}_{-0.04}$ $1.34^{+0.07}_{-0.07}$ $1.12^{+0.03}_{-0.04}$
$\log_{10}M_{\text{BH}}$ $10.50^{+0.10}_{-0.32}$ $10.55^{+0.10}_{-0.07}$ $10.15^{+0.17}_{-0.30}$ $10.33^{+0.07}_{-0.13}$ $10.59^{+0.04}_{-0.10}$
$\epsilon^{\text{sh}}_{1}$ $-0.03^{+0.01}_{-0.01}$ $0.00^{+0.01}_{-0.01}$ $-0.03^{+0.01}_{-0.01}$ $-0.03^{+0.01}_{-0.01}$ $-0.01^{+0.01}_{-0.01}$ $-0.02^{+0.02}_{-0.01}$ $-0.02^{+0.01}_{-0.01}$
$\epsilon^{\text{sh}}_{2}$ $-0.06^{+0.01}_{-0.01}$ $-0.05^{+0.01}_{-0.01}$ $-0.06^{+0.01}_{-0.01}$ $-0.06^{+0.01}_{-0.01}$ $-0.06^{+0.01}_{-0.01}$ $-0.05^{+0.01}_{-0.01}$ $-0.05^{+0.02}_{-0.01}$
$M_{\text{Ein}}$ $5.45^{+0.02}_{-0.03}$ $5.45^{+0.02}_{-0.02}$ $5.44^{+0.02}_{-0.02}$ $5.43^{+0.02}_{-0.02}$ $5.43^{+0.02}_{-0.02}$ $5.46^{+0.02}_{-0.03}$ $5.45^{+0.03}_{-0.03}$